OpenStudy (anonymous):

What is the slope of the line that passes through the points (–2, 5) and (1, 4)?

OpenStudy (chriss):

\[\frac{5-4}{-2-1}\] \[\frac{1}{-3}\] \[-\frac{1}{3}\]